# Indirect forces in disc-planet interaction

### v1

@@ 3.2 High-$M_{\mathrm{p}}$ (nonlinear) regime
In this case, panel (a) shows that there is now a noticeable difference between the density wake shapes in real space obtained with and without the IT; this difference grows with the distance traveled by the wave in the outer disc. Panel (b) shows that in the inner disc the deviations between the calculations with and without the IT, while small, become noticeable for all $m$. The power in the $m=1$ mode in the inner disc is still very low compared to $m\geq 2$ modes, as expected from linear theory. However, it is noticeably non-zero and slowly increases as $R\to 0$, in contradiction with the linear expectations of pure exponential damping. Our tests demonstrate that this slowly increasing inwards $m=1$ mode is caused by the emergence of non-zero eccentricity in the inner disc for $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$, possibly driven by the mechanisms outlined in Teys2016; see also Crida2025 and the discussion of Figure 11 in Appendix C.1. Disc eccentricity is negligible in the $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ run.
@@
We note that these differences are not due to some secondary nonlinear structures, e.g. vortices, that may appear in simulations at high $M_{\mathrm{p}}$ and drive their own density waves. First, our simulations are run for a very short time, $20P_{\mathrm{p}}$, while the vortices take longer to form due to the Rossby Wave Instability (Cimerman2023). Second, vortices would drive density waves not only in the outer regions but also in the inner disc. However, differences between the solid and dashed curves are small in the inner disc. Third, the deviations between the dashed and solid curves in Figure 3c are larger for lower $m$, as expected from a gradual leakage of power from $m=1$ mode to higher $m$ due to wave nonlinearity.
One curious aspect of the Figure 3c is that the decay of $|\delta\Sigma_{m}|$ in the outer disc is no longer monotonic with $R$. Instead, $|\delta\Sigma_{m}|$ curves show oscillatory behavior, going to zero at certain locations, with the frequency of these radial oscillations being higher for higher $m$. In Appendix A we provide an explanation for this behavior and show that it is caused by the azimuthal spreading of the wake profile caused by its nonlinear evolution (GR01; R02; Dong2011b; Duffell2012; Cimerman2021), a process which is much faster in the $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$ case (Fig. 3a) than in the $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ run (Fig. 2a).
@@ 4.1 Excitation torque density
Figure 4: Illustration of the torque density computed (a,b) without the IT and (c) with the IT included. Data for the outer disc ($R>R_{\mathrm{p}}$) are taken from the athena++ simulations for an adiabatic disc with $h_{\mathrm{p}}=0.1$, $p=q=1$ and $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$. In panel (a), only the direct torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ given by the equation (13) is shown; the solid blue curve is $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ computed based on $\Sigma$ from a simulation without the IT, while the dashed orange curve is the direct torque computed using $\Sigma$ from a simulation with the IT properly included. Panel (b) is a vertical zoom of panel (a) illustrating the torque wiggles (Cimerman2024b). In panel (c) the $\Sigma$ data are from the simulation with the IT included, and the different curves show the full torque $\mathrm{d}T/\mathrm{d}R$ (black solid) defined by equation (12), the direct torque $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ (orange dashed, same curve as in panels (a),(b)), and the indirect torque $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ (green dot-dashed). Note the conspicuous oscillations of growing amplitude exhibited by the indirect and total torque densities as $R$ increases in panel (c).
@@
In Figure 4 we show the results of torque calculations performed on the outputs of two $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ simulations — including and excluding the IT in the calculation of the perturbed $\Sigma$. In the former case we also have a freedom of including or excluding $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ in the calculation of the excitation torque density. Torque densities are normalized by the natural reference unit in disc-planet interaction, $F_{J,0}/R_{\mathrm{p}}$, where $F_{J,0}$ is the characteristic one-sided torque (GT80; GR01)
@@
In panel (a) we show in detail the direct torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ computed in the outer disc ($R>R_{\mathrm{p}}$) via equation (13) for both runs. The solid blue curve uses $\Sigma$ from a simulation in which the IT term was absent (‘IT off’), while for the dashed orange curve $\Sigma$ comes from a simulation with the IT term included (‘IT on’). One can see that the two curves closely follow each other around the peak of $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ near the planet, while further out one notices low-amplitude, quasi-periodic, decaying features — the torque wiggles previously studied in Cimerman2024b. In panel (b) we zoom in vertically to better illustrate the torque wiggles and also to highlight how the different ways of computing $\Sigma$ (with and without the IT) impact $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$. One can see that the two curves show small differences reflecting the differences in the perturbation patterns produced by the two simulations, as described in Section 3. However, both of them eventually decay to zero as the distance from the planet increases. So far, the picture is consistent with other existing calculations of the excitation torque density (e.g DAngelo2008; Arza18; Cimerman2024b).
@@
To ensure that these torque oscillations are not an artefact of the nonlinear calculation, in Figure 5 we show $\mathrm{d}T/\mathrm{d}R$ (dot-dashed orange curve) obtained using a linear calculation (see Section 2.4) with the IT fully included in calculations of both $\Sigma$ and $\mathrm{d}T/\mathrm{d}R$ and scaled to $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ used in simulations. For comparison, the solid blue curve is the simulation-based $\mathrm{d}T/\mathrm{d}R$, i.e. same as the black curve in Figure 4c. One can see that the linear calculation reproduces growing oscillations of the torque density with the same periodicity as the fully nonlinear simulations. The difference in the amplitudes of oscillations will be addressed later in Section 4.4. But the clear presence of $\mathrm{d}T/\mathrm{d}R$ oscillations in both linear and nonlinear calculations implies that they are a real physical effect.
The oscillatory behaviour of $\mathrm{d}T/\mathrm{d}R$ in the outer disc may seem puzzling and raises natural questions regarding the angular momentum conservation in the disc. It calls for deeper analysis to understand its origin, which is what we do in the rest of this section.
@@ 4.2 Wave AMF and integrated torque
Torque exerted by a planet on the disc, as described in the previous Section, contributes to the angular momentum of the planet-driven density wave (GT79; GT80). The wave carries this angular momentum away from the planet, a process which is characterized by another key angular momentum characteristic — the wave angular momentum flux (AMF), $F_{J}(R)$. In the astrocentric frame, AMF represents the amount of angular momentum carried by the wave across a cylinder of radius $R$ centered on the star per unit of time. Mathematically, wave AMF can be expressed as
@@
Wave angular momentum conservation (enforced by our thermodynamic assumptions, see Miranda2020I) dictates that in the absence of explicit dissipation, the radial gradient of $F_{J}(R)$ (i.e. the amount of angular momentum added to the wave per unit radius and time) should be equal to the torque density $\mathrm{d}T(R)/\mathrm{d}R$ at the same radius (e.g. Cordwell2024),
@@
Indeed, according to Gold1989, in an ideal (non-dissipative) fluid perturbed by a rigidly rotating external potential, no secular changes in the angular momenta of fluid elements are possible (except at corotation). Therefore, the disc does not change its state and all angular momentum injected by the external potential gets transported away by internal stresses, i.e. waves. Another way to phrase this and interpret equation (17) is to say that $F_{J}(R)$ should match the full (integrated) torque $T(R)$ exerted by the planet on the disc between $R_{\mathrm{p}}$ and $R$,
@@
To set the stage, we start with a linear calculation, akin to Section 2.4. In Figure 6 we show the AMF and integrated torques obtained through the linear framework and normalized by $F_{J,0}$. In panel (a) we illustrate a calculation that neglects the IT altogether, both in the computation of the perturbed $\Sigma$ and in the torque evaluation, in which case $T_{\mathrm{id}}=0$ and $T=T_{\mathrm{d}}$. The solid red curve shows the AMF $F_{J}(R)$ while the dashed orange curve is $T_{\mathrm{d}}(R)=T(R)$. The two curves closely follow each other with just a small constant vertical offset, which can be attributed to the corotation torque effects. Both curves show a rapid rise near the planet, where $\mathrm{d}T/\mathrm{d}R$ peaks, but then flatten out with small, decaying oscillations on top — the torque wiggles (Cimerman2024b) manifesting themselves in both the integrated torque and AMF. This is a standard behaviour previously seen in all calculations ignoring the IT (e.g. Dong2011; RP12).
@@
Figure 6: Radial profiles of the angular momentum flux $F_{J}$ of the planet-driven density wave (solid red), and integrated direct torque $T_{\mathrm{d}}$ (dashed orange), indirect torque $T_{\mathrm{id}}$ (dot-dashed green), and total torque $T$ (solid black). These are linear calculations (see Section 2.4) with no wave dissipation (unlike in Figure 7). In panel (a) IT is fully neglected, $T_{\mathrm{id}}=0$, and $T=T_{\mathrm{d}}$. In panel (b) both the $\Sigma$ perturbation and the torque calculation account for the IT. Except for the small vertical offset (due to the corotation effects), $T(R)$ and $F_{J}(R)$ curves follow each other, which is also illustrated in the inset (panel (c), same colors scheme as in (b) but with the red curve dashed), showing that their radial derivatives are the same, see Eq. (17). In panel (a), the low-amplitude features of decaying amplitude in $T$ and $F_{J}$ are the torque wiggles (Cimerman2024b), whereas the oscillations of increasing amplitude in panel (b) are due to the indirect contribution to the torque. See text for details.
@@
These linear calculations are further corroborated by the results of our nonlinear simulations in Figure 7, where we show $F_{J}(R)$ and integrated torques from two simulations in the astrocentric frame with a $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ planet. In panel (a) we again do not account for the IT both when computing the disc structure and when computing the torque, so that $T(R)=T_{\mathrm{d}}(R)$. One can see that near the planet $F_{J}(R)$ and $T(R)=R_{\mathrm{d}}(R)$ curves again follow each other closely (up to a constant vertical offset), which is a consequence of equation (17) since the wave does not dissipate initially. But then around $1.4R_{\mathrm{p}}$ the density wave shocks (GR01; R02) and starts to dissipate. Its AMF $F_{J}$ starts to decay since wave angular momentum gets transferred to the disc fluid (see Section 4.3). As a result, the $F_{J}$ curve falls below $T_{\mathrm{d}}$ and then steadily decays as $R$ increases; a careful inspection shows the signs of weak torque wiggles (Cimerman2024b) imprinted on both the $T(R)$ and $F_{J}(R)$ curves. This picture is in agreement with a number of earlier studies (Dong2011b; Miranda2019II; Cimerman2021; Cimerman2024b) that did not include the IT.
@@
We can now re-assess the issue of angular momentum conservation when including the IT, as highlighted at the end of Section 4.1. What Figures 6 and 7 show is that accounting for the IT in torque calculations is not an impediment to the angular momentum conservation, but is in fact an integral part. Indeed, if we set $T_{\mathrm{id}}=0$ and $T=T_{\mathrm{d}}$ but kept the IT in the calculation of $\delta\Sigma$, the AMF would still exhibit prominent oscillations in the outer disc. But now, these oscillations would be unmatched by the torque acting on the disc, meaning that the angular momentum of the density wave would not be conserved. Thus, conservation of the wave angular momentum actually requires one to properly include $T_{\mathrm{id}}$ when the IT is active in the disc structure calculation.

### v3

@@ 3.2 High-$M_{\mathrm{p}}$ (nonlinear) regime
In this case, panel (a) shows that there is now a noticeable difference between the density wake shapes in real space obtained with and without the IT; this difference grows with the distance traveled by the wave in the outer disc. Panel (b) shows that in the inner disc the deviations between the calculations with and without the IT, while small, become noticeable for all $m$. The power in the $m=1$ mode in the inner disc is still very low compared to $m\geq 2$ modes, as expected from linear theory. However, it is noticeably non-zero and slowly increases as $R\to 0$, in contradiction with the linear expectations of pure exponential damping. Our tests demonstrate that this slowly increasing inwards $m=1$ mode is caused by the emergence of non-zero eccentricity in the inner disc for $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$, possibly driven by the mechanisms outlined in Teyssandier and Ogilvie (2016); see also Crida et al. (2025) and the discussion of Figure 12 in Appendix C.1. Disc eccentricity is negligible in the $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ run.
@@
We note that these differences are not due to some secondary nonlinear structures, e.g. vortices, that may appear in simulations at high $M_{\mathrm{p}}$ and drive their own density waves. First, our simulations are run for a very short time, $20P_{\mathrm{p}}$, while the vortices take longer to form due to the Rossby Wave Instability (Cimerman and Rafikov, 2023). Second, vortices would drive density waves not only in the outer regions but also in the inner disc. However, differences between the solid and dashed curves are small in the inner disc. Third, the deviations between the dashed and solid curves in Figure 3c are larger for lower $m$, as expected from a gradual leakage of power from $m=1$ mode to higher $m$ due to wave nonlinearity.
One curious aspect of the Figure 3c is that the decay of $|\delta\Sigma_{m}|$ in the outer disc is no longer monotonic with $R$. Instead, $|\delta\Sigma_{m}|$ curves show oscillatory behavior, going to zero at certain locations, with the frequency of these radial oscillations being higher for higher $m$. In Appendix A we provide an explanation for this behavior and show that it is caused by the azimuthal spreading of the wake profile caused by its nonlinear evolution (Goodman and Rafikov, 2001; Rafikov, 2002a; Dong et al., 2011b; Duffell and MacFadyen, 2012; Cimerman and Rafikov, 2021), a process which is much faster in the $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$ case (Fig. 3a) than in the $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ run (Fig. 2a).
@@ 4.1 Excitation torque density
Figure 4: Illustration of the torque density computed (a,b) without the IT and (c) with the IT included. Data for the outer disc ($R>R_{\mathrm{p}}$) are taken from the athena++ simulations for an adiabatic disc with $h_{\mathrm{p}}=0.1$, $p=q=1$ and $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$. In panel (a), only the direct torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ given by the equation (13) is shown; the solid blue curve is $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ computed based on $\Sigma$ from a simulation without the IT, while the dashed orange curve is the direct torque computed using $\Sigma$ from a simulation with the IT properly included. Panel (b) is a vertical zoom of panel (a) illustrating the torque wiggles (Cimerman et al., 2024). In panel (c) the $\Sigma$ data are from the simulation with the IT included, and the different curves show the full torque $\mathrm{d}T/\mathrm{d}R$ (black solid) defined by equation (12), the direct torque $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ (orange dashed, same curve as in panels (a),(b)), and the indirect torque $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ (green dot-dashed). Note the conspicuous oscillations of growing amplitude exhibited by the indirect and total torque densities as $R$ increases in panel (c).
@@
In Figure 4 we show the results of torque calculations performed on the outputs of two $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ simulations — including and excluding the IT in the calculation of the perturbed $\Sigma$. In the former case we also have a freedom of including or excluding $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ in the calculation of the excitation torque density. Torque densities are normalized by the natural reference unit in disc-planet interaction, $F_{J,0}/R_{\mathrm{p}}$, where $F_{J,0}$ is the characteristic one-sided torque (Goldreich and Tremaine, 1980; Goodman and Rafikov, 2001)
@@
In panel (a) we show in detail the direct torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ computed in the outer disc ($R>R_{\mathrm{p}}$) via equation (13) for both runs. The solid blue curve uses $\Sigma$ from a simulation in which the IT term was absent (‘IT off’), while for the dashed orange curve $\Sigma$ comes from a simulation with the IT term included (‘IT on’). One can see that the two curves closely follow each other around the peak of $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ near the planet, while further out one notices low-amplitude, quasi-periodic, decaying features — the torque wiggles previously studied in Cimerman et al. (2024). In panel (b) we zoom in vertically to better illustrate the torque wiggles and also to highlight how the different ways of computing $\Sigma$ (with and without the IT) impact $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$. One can see that the two curves show small differences reflecting the differences in the perturbation patterns produced by the two simulations, as described in Section 3. However, both of them eventually decay to zero as the distance from the planet increases. So far, the picture is consistent with other existing calculations of the excitation torque density (e.g D’Angelo and Lubow, 2008; Arzamasskiy et al., 2018; Cimerman et al., 2024).
@@
To ensure that these torque oscillations are not an artefact of the nonlinear calculation, in Figure 5 we show $\mathrm{d}T/\mathrm{d}R$ (dot-dashed orange curve) obtained using a linear calculation (see Section 2.4) with the IT fully included in calculations of both $\Sigma$ and $\mathrm{d}T/\mathrm{d}R$ and scaled to $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ used in simulations. For comparison, the solid blue curve is the simulation-based $\mathrm{d}T/\mathrm{d}R$, i.e. same as the black curve in Figure 4c. One can see that the linear calculation reproduces growing oscillations of the torque density with the same periodicity as the fully nonlinear simulations. The difference in the amplitudes of oscillations will be addressed later in Section 4.4. But the clear presence of growing $\mathrm{d}T/\mathrm{d}R$ oscillations in both linear and nonlinear calculations implies that they are a real physical effect.
We now show that the radially-diverging behavior of the indirect torque $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ arises because of the reflex motion of the central star. Let us a consider a mass (e.g. a fluid element in the disc) in the barycentric frame of a star-planet system with a finite specific angular momentum ${\bf l}_{\mathrm{b}}={\bf r}\times\dot{\bf r}$, where $\bf r$} is the barycentric position of this mass. In the astrocentric coordinates the position of this mass is ${\bf R}={\bf r}-{\bf r}_{\star}$, where $\bf r$$_{\star}$ is the stellar position in the barycentric frame. Then the specific angular momentum ${\bf l}_{\mathrm{a}}$ of that mass in the astrocentric frame is
Moreover, because of planet-induced stellar motion around the barycenter of the system, $\dot{\bf r}_{\star}$ varies in time and so does ${\bf l}_{\mathrm{a}}$, with an infinitely growing amplitude as $r\to\infty$. This implies that in the astrocentric frame an infinitely large torque must be applied to that distant mass, thereby explaining the divergence of the indirect torque for $r,R\to\infty$ as a consequence of stellar reflex motion. These conclusions are corroborated further in Section 4.5.
The oscillatory behaviour of $\mathrm{d}T/\mathrm{d}R$ in the outer disc may still seem puzzling and raises natural questions regarding the angular momentum conservation in the disc. It calls for deeper analysis to understand its origin, which is what we do in the rest of this section.
@@ 4.2 Wave AMF and integrated torque
Torque exerted by a planet on the disc, as described in the previous Section, contributes to the angular momentum of the planet-driven density wave (Goldreich and Tremaine, 1979, 1980). The wave carries this angular momentum away from the planet, a process which is characterized by another key angular momentum characteristic — the wave angular momentum flux (AMF), $F_{J}(R)$. In the astrocentric frame, AMF represents the amount of angular momentum carried by the wave across a cylinder of radius $R$ centered on the star per unit of time. Mathematically, wave AMF can be expressed as
@@
Wave angular momentum conservation (enforced by our thermodynamic assumptions, see Miranda and Rafikov, 2020b) dictates that in the absence of explicit dissipation, the radial gradient of $F_{J}(R)$ (i.e. the amount of angular momentum added to the wave per unit radius and time) should be equal to the torque density $\mathrm{d}T(R)/\mathrm{d}R$ at the same radius (e.g. Cordwell and Rafikov, 2024),
@@
Indeed, according to Goldreich and Nicholson (1989), in an ideal (non-dissipative) fluid perturbed by a rigidly rotating external potential, no secular changes in the angular momenta of fluid elements are possible (except at corotation). Therefore, the disc does not change its state and all angular momentum injected by the external potential gets transported away by internal stresses, i.e. waves. Another way to phrase this and interpret equation (18) is to say that $F_{J}(R)$ should match the full (integrated) torque $T(R)$ exerted by the planet on the disc between $R_{\mathrm{p}}$ and $R$,
@@
To set the stage, we start with a linear calculation, akin to Section 2.4. In Figure 6 we show the AMF and integrated torques obtained through the linear framework and normalized by $F_{J,0}$. In panel (a) we illustrate a calculation that neglects the IT altogether, both in the computation of the perturbed $\Sigma$ and in the torque evaluation, in which case $T_{\mathrm{id}}=0$ and $T=T_{\mathrm{d}}$. The solid red curve shows the AMF $F_{J}(R)$ while the dashed orange curve is $T_{\mathrm{d}}(R)=T(R)$. The two curves closely follow each other with just a small constant vertical offset, which can be attributed to the corotation torque effects. Both curves show a rapid rise near the planet, where $\mathrm{d}T/\mathrm{d}R$ peaks, but then flatten out with small, decaying oscillations on top — the torque wiggles (Cimerman et al., 2024) manifesting themselves in both the integrated torque and AMF. This is a standard behaviour previously seen in all calculations ignoring the IT (e.g. Dong et al., 2011a; Rafikov and Petrovich, 2012).
@@
Figure 6: Radial profiles of the angular momentum flux $F_{J}$ of the planet-driven density wave (solid red), and integrated direct torque $T_{\mathrm{d}}$ (dashed orange), indirect torque $T_{\mathrm{id}}$ (dot-dashed green), and total torque $T$ (solid black). These are linear adiabatic calculations (see Section 2.4) with no wave dissipation (unlike in Figure 7). In panel (a) IT is fully neglected, $T_{\mathrm{id}}=0$, and $T=T_{\mathrm{d}}$. In panel (b) both the $\Sigma$ perturbation and the torque calculation account for the IT. Except for the small vertical offset (due to the corotation effects), $T(R)$ and $F_{J}(R)$ curves follow each other, which is also illustrated in the inset (panel (c), same colors scheme as in (b) but with the red curve dashed), showing that their radial derivatives are the same, see Eq. (18). In panel (a), the low-amplitude features of decaying amplitude in $T$ and $F_{J}$ are the torque wiggles (Cimerman et al., 2024), whereas the oscillations of increasing amplitude in panel (b) are due to the indirect contribution to the torque. See text for details.
@@
These linear calculations are further corroborated by the results of our nonlinear simulations in Figure 7, where we show $F_{J}(R)$ and integrated torques from two simulations in the astrocentric frame with a $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ planet. In panel (a) we again do not account for the IT both when computing the disc structure and when computing the torque, so that $T(R)=T_{\mathrm{d}}(R)$. One can see that near the planet $F_{J}(R)$ and $T(R)=R_{\mathrm{d}}(R)$ curves again follow each other closely (up to a constant vertical offset), which is a consequence of equation (18) since the wave does not dissipate initially. But then around $1.4R_{\mathrm{p}}$ the density wave shocks (Goodman and Rafikov, 2001; Rafikov, 2002a) and starts to dissipate. Its AMF $F_{J}$ starts to decay since wave angular momentum gets transferred to the disc fluid (see Section 4.3). As a result, the $F_{J}$ curve falls below $T_{\mathrm{d}}$ and then steadily decays as $R$ increases; a careful inspection shows the signs of weak torque wiggles (Cimerman et al., 2024) imprinted on both the $T(R)$ and $F_{J}(R)$ curves. This picture is in agreement with a number of earlier studies (Dong et al., 2011b; Miranda and Rafikov, 2019b; Cimerman and Rafikov, 2021; Cimerman et al., 2024) that did not include the IT.
@@
We can now re-assess the issue of angular momentum conservation when including the IT, as highlighted at the end of Section 4.1. What Figures 6 and 7 show is that accounting for the IT in torque calculations is not an impediment to the angular momentum conservation, but is in fact an integral part. Indeed, if we set $T_{\mathrm{id}}=0$ and $T=T_{\mathrm{d}}$ but kept the IT in the calculation of $\delta\Sigma$, the AMF would still exhibit prominent oscillations in the outer disc. But now, these oscillations would be unmatched by the torque acting on the disc, meaning that the angular momentum of the density wave would not be conserved. Thus, conservation of the wave angular momentum in the astrocentric frame actually requires one to properly include $T_{\mathrm{id}}$ when the IT is active in the disc structure calculation.
Figure 8: Radial profiles of the (a) torque density $\mathrm{d}T/\mathrm{d}r$ and (b) integrated torque $T(r)$ (black dashed) and angular momentum flux $F_{J}(r)$ of the planet-driven density wave as functions of the barycentric distance $r$. These profiles are derived from our barycentric simulations performed with disco and using the same parameters as Figure 7. All curves show evidence of low-amplitude torque wiggles (Cimerman et al., 2024) but no diverging oscillations as in Figure 7b.
We should also emphasize that the concept of $T_{\mathrm{id}}$ and the divergent oscillatory behaviors of $T$ and $F_{J}$ are unique to the non-inertial astrocentric frame. In the inertial, e.g. barycentric, frame indirect forces are absent, obviating $T_{\mathrm{id}}$ altogether and changing the behavior of $T$ and $F_{J}$. We illustrate this using a simulation with the same parameters as in Figure 7 but carried out in the barycentric frame using disco, see Section 2.3. In Figure 8 we show the torque density (panel (a)), integrated torque $T$ and angular momentum flux $F_{J}$ (panel (b)) as functions of the barycentric distance $r$, which is not very different from $R$ since $M_{\mathrm{p}}\ll M_{\star}$. The torque-related quantities are computed including torques exerted by both the planet and the star (both of which are displaced from the barycentric frame’s origin), and $F_{J}$ uses the disc velocity components relative to the system’s barycenter. One can see that, as expected, all these metrics show a regular behavior as $r$ increases: $\mathrm{d}T/\mathrm{d}r\to 0$, $T(r)$ saturates at a constant level, and $F_{J}(r)$ steadily decays as a result of nonlinear wave dissipation. While these curves exhibit the decaying torque wiggles, they do not show the divergent oscillations at large $r$, as typical for the non-inertial frame. Their behavior is qualitatively reminiscent of Figures 4a and 7a, in which only the direct gravitational force was considered, although the perturbation structure and coordinate systems are different.